# Centered Moments of Weighted One-Level Densities of $GL(2)$ $L$-Functions

### v3

@@ 4. Weighted Trace Formula
$\sum_{f\in\mathcal{F}_{k}(1)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)\$$=\ \sum_{f\in\mathcal{F}_{k}(1)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)$
$=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}a_{f}(r)(f)\lambda_{f}(m)$
$=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)r^{\frac{k-1}{2}}\lambda_{f}(r)\lambda_{f}(m).$ (4.7)
@@
$\sum_{f\in\mathcal{F}_{k}(1)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)\$$=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)r^{\frac{k-1}{2}}\lambda_{f}(r)\lambda_{f}(m)$
$=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)r^{\frac{k-1}{2}}\sum_{d\mid(r,m)}\lambda_{f}\left(\frac{rm}{d^{2}}\right)$
$=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)r^{\frac{k-1}{2}}\sum_{d\mid(r,m)}\left(\frac{rm}{d^{2}}\right)^{-\frac{k-1}{2}}a_{f}\left(\frac{rm}{d^{2}}\right).$ (4.8)
@@
$\sum_{f\in\mathcal{F}_{k}(1)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)\$$=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)r^{\frac{k-1}{2}}\sum_{d\mid(r,m)}\left(\frac{rm}{d^{2}}\right)^{-\frac{k-1}{2}}a_{f}\left(\frac{rm}{d^{2}}\right)$
$=\ \sum_{d\mid(r,m)}\left(\frac{m}{d^{2}}\right)^{-\frac{k-1}{2}}\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)a_{f}\left(\frac{rm}{d^{2}}\right).$ (4.9)
@@
$\ \frac{2^{k-1}(2\pi)^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)\sum_{d\mid(r,m)}\left(\frac{m}{d^{2}}\right)^{-\frac{k-1}{2}}\left(\frac{r^{2}m}{d^{2}}\right)^{\frac{k}{2}-1}\chi\left(\frac{r^{2}m}{d^{2}}\right)\sigma\left(\left(r,\frac{rm}{d^{2}}\right)\right)$
$+O\left(\frac{(4\pi r)^{k-1}D^{\frac{k}{2}-\frac{1}{2}}}{(k-2)!}\sum_{d\mid(r,m)}\left(\frac{m}{d^{2}}\right)^{-\frac{k-1}{2}}\left(\frac{rm}{d^{2}}\right)^{k-1}\right)$
$=\$$\ \frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)m^{-\frac{1}{2}}\chi(m)\sum_{d\mid(r,m)}d\ \sigma_{1}\left(\left(r,\frac{rm}{d^{2}}\right)\right)$
$+O\left(m^{\frac{k-1}{2}}\frac{(4\pi r^{2})^{k-1}D^{\frac{k}{2}-\frac{1}{2}}}{(k-2)!}\sum_{d\mid(r,m)}d^{-k+1}\right)$
$=\$$\ \frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)m^{-\frac{1}{2}}\chi(m)\sum_{d\mid(r,m)}d\ \sigma_{1}\left(\left(r,\frac{rm}{d^{2}}\right)\right)$
$+O\left(\frac{m^{\frac{k-1}{2}}V^{k}}{(k-2)!}\right),$ (4.10)
where $V$ is a constant depending on $r$ and $D$. We now recall the following divisor sum identity: for an arithmetic function $f$ and a completely multiplicative function $g$,
@@
Taking $g(n)=n$ and $f(n)=\sigma_{1}((r,n))$ in the above identity allows us to simplify the main term in (4.10), giving that it equals
$\frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)\sigma_{1}(r)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)m^{-\frac{1}{2}}\chi(m)\sigma_{1}((r,m))+O\left(\frac{m^{\frac{k-1}{2}}V^{k}}{(k-2)!}\right).$ (4.12)
This is precisely the formula in the statement of the lemma. ∎
Lemma 4.2.
Suppose $k>2$ and $N>1$. Then, for any positive integer $m\nmid N$, the sum of the $m$th Hecke eigenvalue over $\mathcal{F}_{k}(N)^{\prime}$ is equal to
$\frac{1}{\nu(N)}\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)\$$=\ \frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)\sigma_{1}(r)}{(k-2)!}m^{-\frac{1}{2}}\chi(m)\ \sigma_{1}((r,m))$
$+\ O\left(\frac{m^{\frac{k-1}{2}}V^{k}}{N^{\frac{k}{2}}(k-2)!}\right),$ (4.13)
where the implied constant and $V$ depend only on $r$ and $D$, and $\sigma_{1}$ is the divisor sum function.
@@
Let $m=\prod_{j=1}^{\ell}q_{j}^{m_{j}}\nmid N$, where $q_{1},\ldots,q_{j}$ are the distinct primes dividing $m$. Then, the average (4.1) of the $m$th Hecke eigenvalue over $\mathcal{F}_{k}(N)^{\prime}$ is equal to
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\lambda_{\cdot}\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)\right)\$$=\ \left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)^{-\frac{1}{2}}\left(\prod\nolimits_{j=1}^{\ell}\chi(q_{j})^{m_{j}}\right)\sigma_{1}\left(\left(r,\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)\right)$ (4.14)
$+\ O\left(\frac{\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)^{\frac{k-1}{2}}W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}\right),$ (4.15)
where the implied constant and $W$ depend only on $r$ and $D$.
@@
We prove this statement when $k>2$ and $N>1$. We would like to compute
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(\lambda_{\cdot}(m))\ =\ \frac{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)}{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)}.$ (4.16)
By Lemma 4.2, the denominator (setting $m=1$) equals
$\frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)\sigma_{1}(r)}{(k-2)!}+O\left(\frac{V^{k}}{N^{\frac{k}{2}}(k-2)!}\right)=:M_{D}+E_{D},$ (4.17)
and the numerator equals
$\frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)\sigma_{1}(r)}{(k-2)!}m^{-\frac{1}{2}}\chi(m)\ \sigma_{1}((r,m))+\ O\left(\frac{m^{\frac{k-1}{2}}V^{k}}{N^{\frac{k}{2}}(k-2)!}\right)=:M_{N}+E_{N}.$ (4.18)
@@
$=\ \frac{M_{N}}{M_{D}}+\frac{E_{N}-\frac{M_{N}}{M_{D}}E_{D}}{M_{D}+E_{D}}.$ (4.19)
@@
$\frac{M_{N}}{M_{D}}\$$=m^{-\frac{1}{2}}\chi(m)\ \sigma_{1}((r,m))$
$=\ \left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)^{-\frac{1}{2}}\left(\prod\nolimits_{j=1}^{\ell}\chi(q_{j})^{m_{j}}\right)\sigma_{1}\left(\left(r,\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)\right),$ (4.20)
matching the main term (4.14) in the lemma. We finish by demonstrating that the second term in (4.19) has the desired rate of decay (4.15). For clarity of the analysis, we let
@@
$\frac{E_{N}-\frac{M_{N}}{M_{D}}E_{D}}{M_{D}+E_{D}}\$$\ll\ \frac{\left(m^{\frac{k-1}{2}}-m^{-\frac{1}{2}}\chi(m)\sigma_{1}((r,m))\right)C}{M_{D}+E_{D}}$
@@
$=\ m^{\frac{k-1}{2}}\frac{C/M_{D}}{1+E_{D}/M_{D}}.$ (4.22)
The argument in [JK15, Section 9] gives that
$\frac{E_{D}}{M_{D}}\ \ll\ \frac{C}{M_{D}}\ \ll_{r,D}\ \frac{W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}},$ (4.23)
where $W$ depends only on $r$ and $D$. It directly follows that (4.22) has the claimed decay in the statement of the lemma.
The proof for $N=1$ is similar; the extra factor of $1+\frac{i^{k}\tau(\chi)^{2}}{D}$ in the main term (4.4) will cancel coming from both the numerator and the denominator. ∎
@@ 5. Combinatorial Analysis
To review, we are interested in
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum\nolimits_{p\nmid N}\lambda_{\cdot}(p)\ \widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right],$ (5.1)
@@
$\left.\times\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j},\\ m_{j}\equiv n_{j}(2)\end{subarray}}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\lambda_{\cdot}\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)\right)\right].$ (5.2)
@@
$\sum_{\ell=1}^{t}\sum_{n_{1}+\dots+n_{\ell}=t}$ $\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j}\\ m_{j}\equiv n_{j}(2)\\ 1\leq j\leq\ell\end{subarray}}\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)$
$\times\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell})\\ q_{j}\nmid N\end{subarray}}\left(\prod_{j=1}^{\ell}\widehat{\phi}\left(\frac{\log q_{j}}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi(q_{j})^{m_{j}}\log^{n_{j}}q_{j}}{q_{j}^{(n_{j}+m_{j})/2}\log^{n_{j}}Q_{k,N}}\ \sigma_{1}((r,q_{j}^{m_{j}}))\right);$ (5.3)
@@
$\ll_{r,n,D}\frac{W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell})\\ q_{j}\nmid N\end{subarray}}\left(\prod_{j=1}^{\ell}\widehat{\phi}\left(\frac{\log q_{j}}{\log Q_{k,N}}\right)^{n_{j}}\frac{\log^{n_{j}}q_{j}}{q_{j}^{n_{j}/2}\log^{n_{j}}Q_{k,N}}q_{j}^{\frac{m_{j}(k-1)}{2}}\right).$ (5.4)
Note that we have absorbed the combinatorial sum in (5.2) into the implied constant in (5.4) as it is bounded by a polynomial in $t\leq n$. It is also worth reiterating that $W$ is a constant depending only on $r$ and $D$.
@@
$\ll$ $\ \frac{W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}\prod_{j=1}^{\ell}\left(\sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\log^{n_{j}}q}{q^{n_{j}/2}\log^{n_{j}}Q_{k,N}}q_{j}^{\frac{m_{j}(k-1)}{2}}\right).$ (5.5)
@@
$\ll$ $\ \frac{W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}Q_{k,N}^{\frac{nk\alpha}{2}-n\alpha+n},$ (5.6)
recalling that $m_{j}\leq n_{j}$, $\ell\leq n$, and $\sum n_{j}=n$. Since $Q_{k,N}=k^{2}N$, if $\alpha<\frac{1}{2n}$, then (5.6) vanishes as $k\to\infty$ or $k+N\to\infty$. ∎
Lemma 5.1 licenses us to focus exclusively on the main term contribution (5.3) in our analysis of (5.2). Applying the principle of inclusion-exclusion to (5.3), we swap the sum and the product, reversing the idea in (3.7) and picking up the error term $O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$:
$\sum_{\ell=1}^{t}\sum_{n_{1}+\cdots+n_{\ell}=t}$ $\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j}\\ m_{j}\equiv n_{j}(2)\\ 1\leq j\leq\ell\end{subarray}}\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)$
$\times\prod_{j=1}^{\ell}\left(\sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi(q)^{m_{j}}\log^{n_{j}}q}{q^{(n_{j}+m_{j})/2}\log^{n_{j}}Q_{k,N}}\ \sigma_{1}\left(\left(r,q^{m_{j}}\right)\right)\right).$ (5.7)
We justify this step in tedious detail in Appendix A. We thereby reduce the problem to analyzing
$\sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log R}\right)^{n_{j}}\frac{\chi(q)^{m_{j}}\log^{n}q}{q^{(n+m)/2}\log^{n}R}\ \sigma_{1}\left(\left(r,q^{m}\right)\right)$ (5.8)

### v4

@@ 4. Weighted Trace Formula
(4.9) $\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}(f)\lambda_{f}(m)\$$=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)r^{\frac{k-1}{2}}\lambda_{f}(r)\lambda_{f}(m).$
@@
(4.10) $\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}(f)\lambda_{f}(m)\$$=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)r^{\frac{k-1}{2}}\sum_{d\mid(r,m)}\lambda_{f}\left(\frac{rm}{d^{2}}\right)$
(4.11) $=\ \sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)r^{\frac{k-1}{2}}\sum_{d\mid(r,m)}\left(\frac{rm}{d^{2}}\right)^{-\frac{k-1}{2}}a_{f}\left(\frac{rm}{d^{2}}\right).$
@@
(4.12) $\sum_{f\in\mathcal{F}_{k}(1)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)\$$=\ \sum_{d\mid(r,m)}\left(\frac{m}{d^{2}}\right)^{-\frac{k-1}{2}}\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)a_{f}\left(\frac{rm}{d^{2}}\right).$
@@
$\ \frac{2^{k-1}(2\pi)^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)\sum_{d\mid(r,m)}\left(\frac{m}{d^{2}}\right)^{-\frac{k-1}{2}}\left(\frac{r^{2}m}{d^{2}}\right)^{\frac{k}{2}-1}\chi\left(\frac{r^{2}m}{d^{2}}\right)\sigma_{1}\left(\left(r,\frac{rm}{d^{2}}\right)\right)$
$+O\left(\frac{(4\pi r)^{k}D^{\frac{k}{2}}}{N^{\frac{k}{2}}(k-2)!}\sum_{d\mid(r,m)}\left(\frac{m}{d^{2}}\right)^{-\frac{k}{2}}\left(\frac{rm}{d^{2}}\right)^{k-1}\right)$
$=\$$\ \frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)m^{-\frac{1}{2}}\chi(m)\sum_{d\mid(r,m)}d\cdot\sigma_{1}\left(\left(r,\frac{rm}{d^{2}}\right)\right)$
(4.13) $+O\left(m^{\frac{k}{2}-1}\frac{(4\pi r^{2})^{k}D^{\frac{k}{2}}}{N^{\frac{k}{2}}(k-2)!}\sum_{d\mid(r,m)}d^{-k+1}\right).$
Noting that
(4.14) $\sum_{d\mid(r,m)}d^{-k+1}\leq\sum_{d\mid r}d^{-k+1}\leq\sum_{d=1}^{r}d^{-k+1}\ll_{r}r^{-k},$
we see (4.13) is equal to
$\frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)m^{-\frac{1}{2}}\chi(m)S_{r}(m)$
(4.15) $+O\left(m^{\frac{k}{2}-1}\frac{(4\pi r)^{k-1}D^{\frac{k-1}{2}}}{N^{\frac{k}{2}}(k-2)!}\right).$
which is precisely the claimed formula in the statement of the lemma. ∎
Lemma 4.2.
Suppose $k>2$ and $N>1$. Let $\sigma_{1}$ be the divisor sum function, and $S_{r}(m):=\sum_{d\mid(r,m)}d\cdot\sigma_{1}\left(\left(r,\frac{rm}{d^{2}}\right)\right)$. Then, for any integer $m\geq 1$ with $(m,D)=1$, the sum of the $m$th Hecke eigenvalue over $\mathcal{F}_{k}(N)$ is equal to
$\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}(f)\lambda_{f}(m)$
$=\$$\frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}m^{-\frac{1}{2}}\chi(m)S_{r}(m)$
@@
where the implied constant depends only on $r$ and $D$.
@@
Suppose $k>2$ and $N\geq 1$. Let $\sigma_{1}$ be the divisor sum function, and $S_{r}(m):=\sum_{d\mid(r,m)}d\cdot\sigma_{1}\left(\left(r,\frac{rm}{d^{2}}\right)\right)$. Let $m=\prod_{j=1}^{\ell}q_{j}^{m_{j}}$, where $q_{1},\ldots,q_{j}$ are the distinct primes dividing $m$, and $(m,DN)=1$. Then, the average (4.1) of the $m$th Hecke eigenvalue over $\mathcal{F}_{k}(N)$ is equal to
(4.17) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)}\left(\lambda_{\cdot}\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)\right)\$$=\ \left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)^{-\frac{1}{2}}\left(\prod\nolimits_{j=1}^{\ell}\chi(q_{j})^{m_{j}}\right)S_{r}\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)$
(4.18) $+\ O\left(\frac{\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)^{\frac{k-1}{2}}W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}\right),$
where the implied constant depends only on $r$ and $D$, and $W$ depends only on $D$.
@@
First, we prove this statement when $k>2$ and $N>1$. We would like to compute
(4.19) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)}(\lambda_{\cdot}(m))\ =\ \frac{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)}{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)}.$
By Lemma 4.2 (with $m=1$), the denominator equals
(4.20) $M_{D}+E_{D}\ :=\ \frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}+O\left(\frac{(4\pi r)^{k}D^{\frac{k}{2}}}{N^{\frac{k}{2}}(k-2)!}\right),$
where $M_{D}$ and $E_{D}$ denote the main and error terms, respectively. Following this convention, the numerator equals
(4.21) $M_{N}+E_{N}\ :=\ \frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}m^{-\frac{1}{2}}\chi(m)S_{r}(m)+\ O\left(m^{\frac{k-1}{2}}\frac{(4\pi r)^{k}D^{\frac{k}{2}}}{N^{\frac{k}{2}}(k-2)!}\right).$
@@
(4.22) $=\ \frac{M_{N}}{M_{D}}+\frac{E_{N}-\frac{M_{N}}{M_{D}}E_{D}}{M_{D}+E_{D}}.$
@@
$\frac{M_{N}}{M_{D}}\$$=m^{-\frac{1}{2}}\chi(m)\ S_{r}(m)$
(4.23) $=\ \left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)^{-\frac{1}{2}}\left(\prod\nolimits_{j=1}^{\ell}\chi(q_{j})^{m_{j}}\right)S_{r}\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right),$
matching the main term (4.17) in the lemma. We finish by demonstrating that the second term in (4.22) has the desired rate of decay from (4.18). For clarity of the analysis, we let
@@
$\frac{E_{N}-\frac{M_{N}}{M_{D}}E_{D}}{M_{D}+E_{D}}\$$\ll\ \frac{\left(m^{\frac{k-1}{2}}-m^{-\frac{1}{2}}\chi(m)S_{r}(m)\right)C}{M_{D}+E_{D}}$
@@
(4.25) $=\ m^{\frac{k-1}{2}}\frac{C/M_{D}}{1+E_{D}/M_{D}}.$
Relating $C$ and $M_{D}$ to the variables $C_{0}$ and $F_{0}$ (specialized for $s=\frac{1}{2}$) defined in [KR19, 160-161], we have
(4.26) $M_{D}\ =\ \frac{r^{\frac{k}{2}-1}}{\chi(r)}F_{0},\quad C\ll_{r,D}C_{0}$
In this regards, the argument in [KR19, 161], which originates from [JK15, Section 9], carries through in our analysis:
(4.27) $\frac{E_{D}}{M_{D}}\ \ll_{r,D}\ \frac{C}{M_{D}}\ \ll_{r,D}\frac{C_{0}}{r^{\frac{k}{2}}F_{0}}\ \ll_{r,D}\ \frac{(4\pi De)^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}.$
Taking $W=4\pi De$, it follows that (4.25) has the decay claimed in the statement of the lemma.
The proof for $N=1$ is similar; the extra factor of $\left(1+\frac{i^{k}\tau(\chi)^{2}}{D}\right)$ in the main term (4.6) will cancel coming from both the numerator and the denominator. ∎
@@ 5. Combinatorial Analysis
We are interested in
(5.1) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum\nolimits_{p\nmid N}\lambda_{\cdot}(p)\ \widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right],$
@@
$\sum_{\begin{subarray}{c}1\leq\ell\leq t,\\ n_{1}+\dots+n_{\ell}=t\end{subarray}}\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell}),\\ q_{j}\text{ distinct},\\ q_{j}\nmid N\end{subarray}}$ $\left[\left(\prod_{j=1}^{\ell}\widehat{\phi}\left(\frac{\log q_{j}}{\log Q_{k,N}}\right)^{n_{j}}\frac{\log^{n_{j}}q_{j}}{q_{j}^{n_{j}/2}\log^{n_{j}}Q_{k,N}}\right)\right$.
(5.2) $\left.\times\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j},\\ m_{j}\equiv n_{j}(2)\end{subarray}}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\lambda_{\cdot}\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)\right)\right].$
@@
$\sum_{\ell=1}^{t}\sum_{n_{1}+\dots+n_{\ell}=t}$ $\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j}\\ m_{j}\equiv n_{j}(2)\\ \forall 1\leq j\leq\ell\end{subarray}}\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)$
(5.3) $\times\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell}),\\ q_{j}\text{ distinct},\\ q_{j}\nmid N\end{subarray}}\left(\prod_{j=1}^{\ell}\widehat{\phi}\left(\frac{\log q_{j}}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi(q_{j})^{m_{j}}\log^{n_{j}}q_{j}}{q_{j}^{(n_{j}+m_{j})/2}\log^{n_{j}}Q_{k,N}}\ S_{r}\left(q_{j}^{m_{j}}\right)\right);$
@@
(5.4) $\ll_{r,t,D}\frac{W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell}),\\ q_{j}\text{ distinct},\\ q_{j}\nmid N\end{subarray}}\left(\prod_{j=1}^{\ell}\widehat{\phi}\left(\frac{\log q_{j}}{\log Q_{k,N}}\right)^{n_{j}}\frac{\log^{n_{j}}q_{j}}{q_{j}^{n_{j}/2}\log^{n_{j}}Q_{k,N}}q_{j}^{\frac{m_{j}(k-1)}{2}}\right).$
We highlight that (5.3) uses that $S_{r}$, being a convolution of multiplicative functions, is multiplicative. Also, note that we absorbed the combinatorial sum in (5.2) into the implied constant in (5.4) as it is bounded by $2^{t}$.
@@
(5.5) $\ll$ $\ \frac{W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}\prod_{j=1}^{\ell}\left(\sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\log^{n_{j}}q}{q^{n_{j}/2}\log^{n_{j}}Q_{k,N}}q_{j}^{\frac{m_{j}(k-1)}{2}}\right).$
@@
(5.6) $\ll$ $\ \frac{W^{k}}{N^{\frac{k-1}{2}}k^{\frac{k}{2}-1}}Q_{k,N}^{\frac{nk\alpha}{2}-n\alpha+n},$
recalling that $m_{j}\leq n_{j}$, $\ell\leq t$, and $\sum_{j=1}^{\ell}n_{j}=t$. Since $Q_{k,N}=k^{2}N$, if $\alpha<\frac{1}{2n}$, then (5.6) vanishes as $k\to\infty$ or $k+N\to\infty$. ∎
Now, we turn to analyze the main term contribution (5.3). Applying the principle of inclusion-exclusion to (5.3), we swap the sum and the product, up to an error term of $O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$:
$\sum_{\ell=1}^{t}\sum_{n_{1}+\cdots+n_{\ell}=t}$ $\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j}\\ m_{j}\equiv n_{j}(2)\\ \forall 1\leq j\leq\ell\end{subarray}}\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)$
(5.7) $\times\prod_{j=1}^{\ell}\left(\sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi(q)^{m_{j}}\log^{n_{j}}q}{q^{(n_{j}+m_{j})/2}\log^{n_{j}}Q_{k,N}}\ S_{r}\left(q^{m_{j}}\right)\right).$
We justify this step in detail in Appendix A. We thereby reduce the problem to analyzing
(5.8) $\mathcal{S}(n_{j},m_{j},\phi,\chi)\ :=\ \sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi(q)^{m_{j}}\log^{n_{j}}q}{q^{(n_{j}+m_{j})/2}\log^{n_{j}}Q_{k,N}}\ S_{r}\left(q^{m_{j}}\right)$